OpenStudy (anonymous):

Which equation below passes through the points (1, –1) and (8, 2)? A. y+1 = 3/7x - 24/7 B. 3x – 7y = 10 C. 7x – 3y = 59 D. y+1 = 7/3x -56/3

OpenStudy (anonymous):

Can anybody please help?

OpenStudy (luigi0210):

First find the slope

OpenStudy (texaschic101):

Lets find the slope... slope (m) = (y2 - y1)/(x2 - x1) set 1: (1,-1) set 2: (8,2) m = (2 -(- 1) / (8 - 1) m = 3/7 <-- this is your slope y - y1 = m(x - x1) using either of your points...I will use (1,-1) y - (-1) = 3/7(x - 1) y + 1 = 3/7x - 3/7 y - 3/7x = -3/7 - 1 -3/7x + y = -3/7 - 7/7 -3/7x + y = -10/7 (multiply by 7) -3x + 7y = -10 (multiply by -1 to make x positive) 3x - 7y = 10

OpenStudy (anonymous):

Thanks a lot!!! :)

OpenStudy (texaschic101):

no problem...glad to help :)